Joan is 7 years older than twice her son Rommel. Five years ago she was 3 years less than thrice as old as her son. Find the present age of Rommel.
Answer by JulietG(1812)

J = 2R + 7
J-5 = 3(R-5) - 3
Distribute
J-5 = 3R-15-3
J-5 = 3R-18
Add 5 to each side
J = 3R - 13
Take that value of J and substitute it for J in the first equation
J = 2R + 7
3R-13 = 2R+7
Subtract 2R from each side
R - 13 = 7
Add 13 to each side
R = 20
If Rommel is 20, then Joan is 47.

Prove it:
Five years ago (47-5=42) she was 3 years less than thrice as old as her son 20-5=15).
42 = (3*15)-3
42=45-3
42=42